OpenStudy (anonymous):

Hey y'all! I need help on word problems! Part A: Max rented a motorbike at $465 for 5 days. If he rents the same motorbike for a week, he has to pay a total rent of $625. Write an equation in the standard form to represent the total rent (y) that Max has to pay for renting the motorbike for x days. (4 points) Part B: Write the equation obtained in Part A using function notation. (2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals.

OpenStudy (anonymous):

Part A: Well, each day it costs $93 to rent. So, y= x*93.

OpenStudy (anonymous):

A.625-465=$160 he has to pay for 2 days=160 $ rent paid for 1 day=260/2=$80 amount paid for 5days=8*5=$400 But he paid $465 Difference 465-400=$65 so every one has to pay $65+rent for the no. of days total rent for x days y=65+80 x (dollars) B. f(x)=65+80 x